# Emotional Speech-driven 3D Body Animation via Disentangled Latent Diffusion

### v2

@@ 3 Method
The AMUSE pipeline consists of two separately trained networks. The audio disentanglement module, which encodes input speech into latent vectors for content, emotion, and style is described in Sec. 3.2. The main architecture is described in Sec. 3.3. It consists of a 3D human motion prior coupled with a latent diffusion model. It takes random noise (or partially denoised latent vectors) on the input and outputs a human motion sequence. We introduce broader applications in gesture editing in Sec. 3.4.
@@ 3.2 Speech Disentanglement Model
The audio module is trained with a multiple loss terms that ensure that the three latent spaces are properly disentangled. In addition to the standard autoencoder reconstruction loss, we also employ three cross-reconstruction losses, in which we enforce the correct reconstruction of the audio signal where we modify one of the content, style or emotion latents. Additionally, we employ three loss terms on the latent vector predictions – namely emotion and style classification losses over $e$ and $s$, and a content similarity loss between pairs of two content latent vectors extracted from audios that have the same spoken content. For a detailed description of the encoder–decoder architecture, a formal definition of the loss functions and a detailed description of the training process please refer to the Appendix
@@ 3.3 Gesture Generation Model
Similar to [15, 80], our motion prior network is a VAE transformer architecture with encoder $\mathcal{P}_{E}$ and decoder $\mathcal{P}_{D}$. Specifically, both $\mathcal{P}_{E}$ and $\mathcal{P}_{D}$ follow a U-Net-like [89] structure with skip connections between transformer blocks (see Appendix for details). The positional embeddings are learnable and injected into each multi-head attention layer, following the design of Carion et al. [13]. Formally, the encoder takes a sequence of $T$ frames of the SMPL-X pose vectors $m^{1:T}\in\mathbb{R}^{6J\times T}$ and the first two tokens of its output, $\mu\in\mathbb{R}^{d_{m}}$ and $\Sigma\in\mathbb{R}^{d_{m}\times d_{m}}$ are used to extract the motion latent $z_{m}\in\mathbb{R}^{d_{m}}$ via the reparametrization trick. The decoder takes zero positional encodings as query input and the motion latent is fed as memory to every cross-attention transformer layer, producing the reconstructed motion $\hat{m}^{1:T}$.
@@
We employ DDIM [94] to infer high quality conditional motion samples with a small number of denoising timesteps. During inference we draw a sample vector from $\mathcal{N}(\textbf{0},\textbf{I})$ to iteratively denoise in reversed timesteps. The denoised sample is then passed through the decoder $\mathcal{P}_{D}(z_{\tilde{m}^{1:T}})$ to obtain motion $\tilde{m}^{1:T}$
@@ 4 Implementation Details
The BEAT [58] mocap sequences, captured in a Vicon system at 120 Hz, are downsampled to 30 Hz and processed using MoSh++ [61, 65] to obtain SMPL-X parameters. Given a sequence of 3D mocap marker positions, we jointly optimize SMPL-X shape and pose parameters, 3D body translation, and embedding of the mocap markers in the SMPL-X surface. Once processed, the sequences are then divided according to the emotion annotations in the BEAT dataset. We use sequences of English speaking subjects in monologue speaking style for training and evaluating AMUSE. For each sequence we draw $m^{1:L}$ at 30 FPS and concatenate with audio content $c$, emotion $e$, and style $s$ latent vectors. Then, we segment it to 10-sec windows $T$, beginning from the timestamp 0 and discarding additional unaligned information at the end. This preprocessing choice allows us to train transformer networks without masking. We provide additional data processing information in the Appendix
@@
The motion prior is a VAE encoder–decoder with 9 layers and 4 heads, following Chen et al. [15]. The encoder–decoder is a U-Net-like transformer with residual connections. Learnable positional embeddings are injected in each multi-head attention layer. We have a linear projection at the start and the end of our motion prior network. The KL divergence term is weighted with a factor of $41e-$ 4.
@@ 5 Experiments
We evaluate the performance of the speech disentanglement model quantitatively using classification accuracy and F1 scores on emotion and style. The accuracy is computed as average scores for all 8 emotion as well style categories that are part of the test dataset. The emotion and style accuracy is 91.53% and 96.06%, respectively. The emotion F1 score and style F1 scores are 0.914 and 0.960, respectively. See the Appendix for ablations and a detailed metric analysis.
@@
We evaluate the performance of our gesture generation model quantitatively, qualitatively, and perceptually against following methods: TalkSHOW [107] and the re-implementation of Habibie et al. [34] provided by the TalkSHOW authors in the official TalkSHOW release [106], DiffuseStyleGesture (DSG) [103], MoGlow [37], and CaMN [58]. Additionally, we adapt TalkSHOW to include categorical emotion labels as input along with the existing architecture that only allows one-hot encodings of personal style. We then retrain it on our training data. We refer to it as TalkSHOW-BEAT. There are some concurrent works [7, 5, 59], which introduce methods for gesture generation from speech, however, direct comparison is hindered by the unavailability of released code our task. Refer to the Appendix for the ablation experiments, the emotion and style editing experiments, and their quantitative evaluation.
@@ 5.1 Quantitative Evaluation
Table 1: Gesture quantitative results. We compare our methods against several SOTA methods using metrics explained in Sec. 5.1. We observe that AMUSE outperforms in all scores compared to baseline methods. Additionally, AMUSE-EmoEdit outperforms in Beat Align, Diversity, and Gesture Emotion Accuracy scores compared to the baseline methods.
To quantitatively evaluate our method’s gesture generations and edited gesture generations, we train a transformer-based encoder architecture (denoted as $M$) similar to Petrovich et al. [80] in an autoencoder setting, where we append a CLS token at the beginning of the motion sequence. $M$ is trained with a cross-entropy emotion classification objective applied to the output CLS token. We train $M$ on the BEAT training dataset and use its features to compute the following metrics: (1) Fréchet gesture distance (FGD): We follow [109, 110, 93] to compute the feature distance between generated and ground truth motion features. (2) Gesture diversity (Div): Similarly to Chen et al. [15], we compute variance across generated features. (3) Gesture emotion accuracy (GA): We report top-1 emotion classification accuracy predicted by a classifier trained on the motion $M$-predicted latents. (4) Beat align (BA): We follow [54, 58], to evaluate the motion-speech correlation in terms of the similarity between the kinematic motion beats and speech audio beats. The kinematic motion beats are directly computed from the generated motion sequences. (5) Semantic-Relevant Gesture Recall (SRGR): We follow Liu et al. [58], to evaluate the semantic relevancy of gestures with GT motion. We use the ground truth semantic scores to compute this metric. The scores are obtained from the BEAT authors, representing a continuous score on a scale 0-1 per gesture style for 4 gesture semantic categories: beat, deictic, iconic, and metaphoric. While comparing with methods that output coarse skeletal data (DSG [103], MoGlow [37], and CaMN [58]), we convert the skeleton motion data into the SMPL-X axis angle representation. For details on the architectures and training of $M$, and the losses, please refer to the Appendix
@@ 5.2 Qualitative Evaluation
In Fig. 4, we demonstrate comparison with baseline methods that output a 3D body mesh: Habibie et al. [34], TalkSHOW [107], TalkSHOW-BEAT, and the BEAT ground truth (GT) [58]. We observe that AMUSE generates gestures that are semantically closer to the speech content and produces expressive emotional gestures closer to the perceived emotion. For example, the GT motion exhibits anger when saying “put them into detention”. AMUSE demonstrates tense posture and aggressive movements comparable with the ground truth data and accurate synchronization with the spoken words. TalkSHOW [107] and Habibie et al. [34] exhibit limited movement and display inferior and static gestures on test audios as seen in the last two rows of Fig. 4. TalkSHOW-BEAT slightly outperforms other baseline methods by demonstrating enhanced synchronized gestures, but it still does not perform as well as AMUSE.
@@
In Fig. 5, our probabilistic model can generate diverse gestures for same input audio.
@@
In Fig. 3 AMUSE demonstrates strong correlation with the spoken utterances as well as different emotions. We observe that our model is able to correlate semantic words to associated gestures. For example, gestures demonstrate forceful actions and tense stance with angry audio “normally get angry” whereas it generates lowered and calm hand positions for sad audio “I went to collect”. Similarly, our generations show hands that are closer to body for fearful audio “in a dangerous situation” while widely open expressing astonishment for happy and surprised audio “people pass by” and “Wow! You are here”.
Refer to captionRefer to caption
@@
We use two audio streams of a female subject for neutral and sad emotion. This experiment edits the subject’s gesture style from moderately controlled hand movements to a sad style with lethargic posture conveying a sense of heaviness, as seen in Fig. 7 (top).
@@
We use audio streams of two male subjects for the happy (ID - 13) and angry (ID - 2) emotion. With the emotion, style and content latent fusion mechanism from two driving audio streams, AMUSE is able to adapt the male (ID - 13) subject’s body gestures from being close to their body to more open with squared tightened shoulders, expressing a shift from happy to angry emotions of a different subject (ID - 2), as shown in Fig. 7 (bottom). Please refer to the supplemental video for qualitative results and comparisons to additional gesture generation methods [58, 4, 116] trained on coarse skeletal data.
@@ Results.
The results of the study are shown in Fig. 6. AMUSE outperforms all competing methods by a considerable margin on both tasks, suggesting that AMUSE’s generations are more appropriate for both the content of the input speech and its emotion compared to the baselines. However, it must be noted that there is still a significant gap between AMUSE and the GT. Please refer to the Appendix for details about the perceptual study.
@@ 6 Conclusion
Acknowledgments. We thank A Cseke, T McConnell and T Alexiadis for help with the perceptual study, B Pellkofer, J Piles-Contreras, and E Fritzler for IT support, and P Kulits, M Petrovich, S Kokane, and S Zojaji for proofreading and insightful discussions. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 860768.
@@
This supplementary material summarizes the video content in Appendix A and provides additional technical details of the speech disentangled model and the gesture generation model in Appendix B and Appendix C, respectively. We provide details about motion extractor model in Appendix D, discussions on the gesture emotion and semantics in Appendix E, details on the data preparation process in Appendix F, a review of state of the art methods in Appendix G, and additional information about the perceptual study in Appendix H.
@@ Appendix B Speech Disentanglement Model
We explain the overall architecture in Sec. B.1 and the encoder–transformer architecture in Sec. B.2. We demonstrate the reconstruction mechanism to enforce disentanglement in Sec. B.3. Finally, we explain the training procedure and loss terms in Sec. B.4.
@@ B.1 Architecture
We illustrate speech disentanglement model architecture in Fig. A.1. The training is conducted over audio of the same utterances spoken under different emotions or spoken by different speakers. Our model consists of three transformer encoders, a transformer fusion, and a transformer decoder. The input filterbank is simultaneously passed through content $E_{c}$, style $E_{s}$, and emotion $E_{e}$ transformer encoders, producing three disentangled latents: content $c$, style $s$, and emotion $e$. The fusion and decoder are transformer-based layers. The transformer–fusion creates a single embedding by applying cross attention on the input triplet embeddings $(c,e,s)$. Finally, the transformer–decoder reconstructs the original filter bank from the compressed single latent embedding produced by the transformer fusion.
Figure A.1: Speech disentanglement model. An input filterbank is given to the three encoders, producing three disentangled latents, which are decoded into a reconstructed filterbank. We here show disentanglement reconstruction for one audio only, please refer Sec. B.3 for its detailed explanation.
@@ B.2 Encoder Transformers
Similar to [27, 25, 74, 31, 28, 26] we employ transfer learning of vision task to our audio task by using pretrained weights of DeiT [98] (88M params) transformer that is fine-tuned on 384x384 images from ImageNet-1k [90]. We use a pretrained DeiT encoder as a component of each of the encoders, as illustrated in Fig. A.3. We linearly embed patches to features embedding of size 768 and feed them into DeiT along with trainable positional embedding of same size (768). We append class token [CLS] and distillation token [DIST] obtained from DeiT at the beginning of each filter bank sequence. We then average the 3-channel inputs of DeiT to obtain a single filterbank channel input. Finally, we use the output of the last DeiT encoder layer and project to 1D latent vector of 256 dimensions each, as our content, emotion, and style latents. We average the [CLS] and [DIST] tokens from DeiT and use it for audio emotion as well as audio style classification tasks for 8 and 30 category labels respectively.
@@ B.3 Reconstruction Concatenations
Fig. A.2 demonstrates a detailed information of the cross-reconstruction mechanism to enforce the audio content, emotion, and style disentanglement. Each audio in the quadruple is encoded and decoded to produce the reconstructed audio filterbank. To enforce content disentanglement, we swap the content latent vectors between different-subject same-emotion audio pairs with same utterances. Similarly, we swap emotion and style latents using audio pairs from the same subject. Specifically, we swap emotion latent vector and style latent vector between same-subject same-emotion audio pairs with different utterances. The procedure is repeated for each audio in the audio quadruples.

### v3

@@ 3 Method
The AMUSE pipeline consists of two separately trained networks. The audio disentanglement module, which encodes input speech into latent vectors for content, emotion, and style is described in Sec.˜3.2. The main architecture is described in Sec.˜3.3. It consists of a 3D human motion prior coupled with a latent diffusion model. It takes random noise (or partially denoised latent vectors) on the input and outputs a human motion sequence. We introduce broader applications in gesture editing in Sec.˜3.4.
@@ 3.2 Speech Disentanglement Model
The audio module is trained with a multiple loss terms that ensure that the three latent spaces are properly disentangled. In addition to the standard autoencoder reconstruction loss, we also employ three cross-reconstruction losses, in which we enforce the correct reconstruction of the audio signal where we modify one of the content, style or emotion latents. Additionally, we employ three loss terms on the latent vector predictions – namely emotion and style classification losses over $e$ and $s$, and a content similarity loss between pairs of two content latent vectors extracted from audios that have the same spoken content. For a detailed description of the encoder–decoder architecture, a formal definition of the loss functions and a detailed description of the training process please refer to the Sup. Mat.
@@ 3.3 Gesture Generation Model
Similar to [15, 80], our motion prior network is a VAE transformer architecture with encoder $\mathcal{P}_{E}$ and decoder $\mathcal{P}_{D}$. Specifically, both $\mathcal{P}_{E}$ and $\mathcal{P}_{D}$ follow a U-Net-like [89] structure with skip connections between transformer blocks (see Sup. Mat. for details). The positional embeddings are learnable and injected into each multi-head attention layer, following the design of Carion et al. [13]. Formally, the encoder takes a sequence of $T$ frames of the SMPL-X pose vectors $m^{1:T}\in\mathbb{R}^{6J\times T}$ and the first two tokens of its output, $\mu\in\mathbb{R}^{d_{m}}$ and $\Sigma\in\mathbb{R}^{d_{m}\times d_{m}}$ are used to extract the motion latent $z_{m}\in\mathbb{R}^{d_{m}}$ via the reparametrization trick. The decoder takes zero positional encodings as query input and the motion latent is fed as memory to every cross-attention transformer layer, producing the reconstructed motion $\hat{m}^{1:T}$.
@@
We employ DDIM [94] to infer high quality conditional motion samples with a small number of denoising timesteps. During inference we draw a sample vector from $\mathcal{N}(\textbf{0},\textbf{I})$ to iteratively denoise in reversed timesteps. The denoised sample is then passed through the decoder $\mathcal{P}_{D}(z_{\tilde{m}^{1:T}})$ to obtain motion $\tilde{m}^{1:T}$ .
@@ 4 Implementation Details
The BEAT [58] mocap sequences, captured in a Vicon system at 120 Hz, are downsampled to 30 Hz and processed using MoSh++ [61, 65] to obtain SMPL-X parameters. Given a sequence of 3D mocap marker positions, we jointly optimize SMPL-X shape and pose parameters, 3D body translation, and embedding of the mocap markers in the SMPL-X surface. Once processed, the sequences are then divided according to the emotion annotations in the BEAT dataset. We use sequences of English speaking subjects in monologue speaking style for training and evaluating AMUSE. For each sequence we draw $m^{1:L}$ at 30 FPS and concatenate with audio content $c$, emotion $e$, and style $s$ latent vectors. Then, we segment it to 10-sec windows $T$, beginning from the timestamp 0 and discarding additional unaligned information at the end. This preprocessing choice allows us to train transformer networks without masking. We provide additional data processing information in the Sup. Mat.
@@
The motion prior is a VAE encoder–decoder with 9 layers and 4 heads, following Chen et al. [15]. The encoder–decoder is a U-Net-like transformer with residual connections. Learnable positional embeddings are injected in each multi-head attention layer. We have a linear projection at the start and the end of our motion prior network. The KL divergence term is weighted with a factor of 1e41e-4.
@@ 5 Experiments
We evaluate the performance of the speech disentanglement model quantitatively using classification accuracy and F1 scores on emotion and style. The accuracy is computed as average scores for all 8 emotion as well style categories that are part of the test dataset. The emotion and style accuracy is 91.53% and 96.06%, respectively. The emotion F1 score and style F1 scores are 0.914 and 0.960, respectively. See the Sup. Mat. for ablations and a detailed metric analysis.
@@
We evaluate the performance of our gesture generation model quantitatively, qualitatively, and perceptually against following methods: TalkSHOW [107] and the re-implementation of Habibie et al. [34] provided by the TalkSHOW authors in the official TalkSHOW release [106], DiffuseStyleGesture (DSG) [103], MoGlow [37], and CaMN [58]. Additionally, we adapt TalkSHOW to include categorical emotion labels as input along with the existing architecture that only allows one-hot encodings of personal style. We then retrain it on our training data. We refer to it as TalkSHOW-BEAT. There are some concurrent works [7, 5, 59], which introduce methods for gesture generation from speech, however, direct comparison is hindered by the unavailability of released code our task. Refer to the Sup. Mat. for the ablation experiments, the emotion and style editing experiments, and their quantitative evaluation.
@@ 5.1 Quantitative Evaluation
Table 1: Gesture quantitative results. We compare our methods against several SOTA methods using metrics explained in Sec.˜5.1. We observe that AMUSE outperforms in all scores compared to baseline methods. Additionally, AMUSE-EmoEdit outperforms in Beat Align, Diversity, and Gesture Emotion Accuracy scores compared to the baseline methods.
To quantitatively evaluate our method’s gesture generations and edited gesture generations, we train a transformer-based encoder architecture (denoted as $M$) similar to Petrovich et al. [80] in an autoencoder setting, where we append a CLS token at the beginning of the motion sequence. $M$ is trained with a cross-entropy emotion classification objective applied to the output CLS token. We train $M$ on the BEAT training dataset and use its features to compute the following metrics: (1) Fréchet gesture distance (FGD): We follow [109, 110, 93] to compute the feature distance between generated and ground truth motion features. (2) Gesture diversity (Div): Similarly to Chen et al. [15], we compute variance across generated features. (3) Gesture emotion accuracy (GA): We report top-1 emotion classification accuracy predicted by a classifier trained on the motion $M$-predicted latents. (4) Beat align (BA): We follow [54, 58], to evaluate the motion-speech correlation in terms of the similarity between the kinematic motion beats and speech audio beats. The kinematic motion beats are directly computed from the generated motion sequences. (5) Semantic-Relevant Gesture Recall (SRGR): We follow Liu et al. [58], to evaluate the semantic relevancy of gestures with GT motion. We use the ground truth semantic scores to compute this metric. The scores are obtained from the BEAT authors, representing a continuous score on a scale 0-1 per gesture style for 4 gesture semantic categories: beat, deictic, iconic, and metaphoric. While comparing with methods that output coarse skeletal data (DSG [103], MoGlow [37], and CaMN [58]), we convert the skeleton motion data into the SMPL-X axis angle representation. For details on the architectures and training of $M$, and the losses, please refer to the Sup. Mat.
@@ 5.2 Qualitative Evaluation
In Fig.˜4, we demonstrate comparison with baseline methods that output a 3D body mesh: Habibie et al. [34], TalkSHOW [107], TalkSHOW-BEAT, and the BEAT ground truth (GT) [58]. We observe that AMUSE generates gestures that are semantically closer to the speech content and produces expressive emotional gestures closer to the perceived emotion. For example, the GT motion exhibits anger when saying “put them into detention”. AMUSE demonstrates tense posture and aggressive movements comparable with the ground truth data and accurate synchronization with the spoken words. TalkSHOW [107] and Habibie et al. [34] exhibit limited movement and display inferior and static gestures on test audios as seen in the last two rows of Fig.˜4. TalkSHOW-BEAT slightly outperforms other baseline methods by demonstrating enhanced synchronized gestures, but it still does not perform as well as AMUSE.
@@
In Fig.˜5, our probabilistic model can generate diverse gestures for same input audio.
@@
In Fig.˜3 AMUSE demonstrates strong correlation with the spoken utterances as well as different emotions. We observe that our model is able to correlate semantic words to associated gestures. For example, gestures demonstrate forceful actions and tense stance with angry audio “normally get angry” whereas it generates lowered and calm hand positions for sad audio “I went to collect”. Similarly, our generations show hands that are closer to body for fearful audio “in a dangerous situation” while widely open expressing astonishment for happy and surprised audio “people pass by” and “Wow! You are here”.
@@
We use two audio streams of a female subject for neutral and sad emotion. This experiment edits the subject’s gesture style from moderately controlled hand movements to a sad style with lethargic posture conveying a sense of heaviness, as seen in Fig.˜7 (top).
@@
We use audio streams of two male subjects for the happy (ID - 13) and angry (ID - 2) emotion. With the emotion, style and content latent fusion mechanism from two driving audio streams, AMUSE is able to adapt the male (ID - 13) subject’s body gestures from being close to their body to more open with squared tightened shoulders, expressing a shift from happy to angry emotions of a different subject (ID - 2), as shown in Fig.˜7 (bottom). Please refer to the supplemental video for qualitative results and comparisons to additional gesture generation methods [58, 4, 116] trained on coarse skeletal data.
@@ Results.
The results of the study are shown in Fig. 6. AMUSE outperforms all competing methods by a considerable margin on both tasks, suggesting that AMUSE’s generations are more appropriate for both the content of the input speech and its emotion compared to the baselines. However, it must be noted that there is still a significant gap between AMUSE and the GT. Please refer to the Sup. Mat. for details about the perceptual study.
@@ 6 Conclusion
Acknowledgments. We thank A Cseke, T McConnell and T Alexiadis for help with the perceptual study, B Pellkofer, J Piles-Contreras, and E Fritzler for IT support, and P Kulits, M Petrovich, S Kokane, and S Zojaji for proofreading and insightful discussions. This work was supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 860768 and by the Swedish Research Council through grant 2020-05187.
@@
This supplementary material summarizes the video content in Appendix˜A and provides additional technical details of the speech disentangled model and the gesture generation model in Appendix˜B and Appendix˜C, respectively. We provide details about motion extractor model in Appendix˜D, discussions on the gesture emotion and semantics in Appendix˜E, details on the data preparation process in Appendix˜F, a review of state of the art methods in Appendix˜G, and additional information about the perceptual study in Appendix˜H.
@@ Appendix B Speech Disentanglement Model
We explain the overall architecture in Sec.˜B.1 and the encoder–transformer architecture in Sec.˜B.2. We demonstrate the reconstruction mechanism to enforce disentanglement in Sec.˜B.3. Finally, we explain the training procedure and loss terms in Sec.˜B.4.
@@ B.1 Architecture
We illustrate speech disentanglement model architecture in Fig.˜A.1. The training is conducted over audio of the same utterances spoken under different emotions or spoken by different speakers. Our model consists of three transformer encoders, a transformer fusion, and a transformer decoder. The input filterbank is simultaneously passed through content $E_{c}$, style $E_{s}$, and emotion $E_{e}$ transformer encoders, producing three disentangled latents: content $c$, style $s$, and emotion $e$. The fusion and decoder are transformer-based layers. The transformer–fusion creates a single embedding by applying cross attention on the input triplet embeddings $(c,e,s)$. Finally, the transformer–decoder reconstructs the original filter bank from the compressed single latent embedding produced by the transformer fusion.
Figure A.1: Speech disentanglement model. An input filterbank is given to the three encoders, producing three disentangled latents, which are decoded into a reconstructed filterbank. We here show disentanglement reconstruction for one audio only, please refer Sec.˜B.3 for its detailed explanation.
@@ B.2 Encoder Transformers
Similar to [27, 25, 74, 31, 28, 26] we employ transfer learning of vision task to our audio task by using pretrained weights of DeiT [98] (88M params) transformer that is fine-tuned on 384x384 images from ImageNet-1k [90]. We use a pretrained DeiT encoder as a component of each of the encoders, as illustrated in Fig.˜A.3. We linearly embed patches to features embedding of size 768 and feed them into DeiT along with trainable positional embedding of same size (768). We append class token [CLS] and distillation token [DIST] obtained from DeiT at the beginning of each filter bank sequence. We then average the 3-channel inputs of DeiT to obtain a single filterbank channel input. Finally, we use the output of the last DeiT encoder layer and project to 1D latent vector of 256 dimensions each, as our content, emotion, and style latents. We average the [CLS] and [DIST] tokens from DeiT and use it for audio emotion as well as audio style classification tasks for 8 and 30 category labels respectively.
@@ B.3 Reconstruction Concatenations
Fig.˜A.2 demonstrates a detailed information of the cross-reconstruction mechanism to enforce the audio content, emotion, and style disentanglement. Each audio in the quadruple is encoded and decoded to produce the reconstructed audio filterbank. To enforce content disentanglement, we swap the content latent vectors between different-subject same-emotion audio pairs with same utterances. Similarly, we swap emotion and style latents using audio pairs from the same subject. Specifically, we swap emotion latent vector and style latent vector between same-subject same-emotion audio pairs with different utterances. The procedure is repeated for each audio in the audio quadruples.